OpenStudy (anonymous):

Help Please? Write an expression to describe a rule for the sequence. Then find the 100th term in the sequence. 5. –4, 1, 6, 11, 16, . . . (1 point) a. –10n + 6; –994 b. 5n – 9; 491 c. 6n – 10; 590 d. –9n + 5; –895 6. 3, 7, 11, 15, 19, 23, . . . (1 point) a. 2n + 1; 201 b. n + 4; 104 c. n + 4; 403 d. 4n – 1; 399

OpenStudy (anonymous):

well... regarding to the first one: as much as I see it is arithmetic sequence: use: \[\LARGE a_n=a_1+(n-1)d\] you have: \[\LARGE d=a_n-a_{n-1}\] \[\LARGE d=a_2-a_1\] \[\LARGE d=1-(-4)\] \[\LARGE d=5\] so: \[\LARGE a_{100}=a_1+(100-1)d\] \[\LARGE a_{100}=-4+99\cdot 5 \] \[\LARGE a_{100}=-4+495 \] \[\LARGE a_{100}=491 \] on the second sequence you have arithmetic sequence too... use these steps , and there you go :)